OpenStudy (anonymous):

if csc theta=-5/3 and theta has its terminal side in quadrant 3, find the exact value of tan 2 theta

OpenStudy (anonymous):

\[\tan 2x=\frac{ 2tanx }{ 1-\tan^2x }\]

OpenStudy (anonymous):

\[sinx=\frac{ -3 }{ 5 }\]

OpenStudy (anonymous):

now get the value of tan x and substitute in my formula

OpenStudy (anonymous):

@mustafa2014 \[\csc x=\frac{ 1 }{ \sin x }\]

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

how do you find the value of tan x?

OpenStudy (anonymous):

\[tanx=\frac{ \sin x }{ \cos x }\]

OpenStudy (anonymous):

\[\cos x=\frac{ -4 }{ 5 }\]

OpenStudy (anonymous):

so tan x= 3/4 right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

so 279.5º?

OpenStudy (anonymous):

the angle should be between 180 and 270

OpenStudy (anonymous):

217 degrees approx.

OpenStudy (anonymous):

I must of done my math wrong then. I plugged in 3/4. so 2tan(3/4)/1-tan^2(3/4) which was 1.8/0.36= 4.87* 180/π. what did i do wrong?

OpenStudy (anonymous):

tan 2x=24/7

OpenStudy (anonymous):

just substitute the value of tanx=3/4 in the formula i gave in the beginning!

OpenStudy (anonymous):

oh ok, I see. Thank you for your help again!